1. Su the following to 2 decimal places. nE(X) = 2, Var(X) = 9, E(Y) =0, Var(Y) = 4, and Corr(X,Y) = 0.25. Find a. Var(X +Y) b. Cov(X, X +Y) c. Corr(X+Y,x-Y).
Answer #1

a)Cov(X,Y)=Corr(X,Y)*sqrt(Var(X)*Var(Y))=0.25*sqrt(9*4)=1.5

Var(X+Y)=Var(X)+Var(Y)+2*Cov(X,Y) =9+4+2*1.5=16

b)

here E(X(X+Y))=E(X2+XY)=E(X2)+E(XY)=(Var(X)+(E(X))2) +(Cov(X,Y)+E(X)*E(Y))=(9+22)+(1.5+2*0)

=13+1.5=14.5

and E(X)*E(X+Y)=2*(2+0)=4

hence Cov(X,X+Y))=E(X*(X+Y))-E(X)*E(X+Y)=14.5-4=10.5

c)

here Var(X+Y)=Var(X)+Var(Y)+2*Cov(X,Y) =9+4+2*1.5=16

Var(X-Y)=Var(X)+Var(Y)-2*Cov(X,Y) =9+4-2*1.5=10

E((X+Y)*(X-Y))=E(X2-Y2) =(Var(X)+(E(X))2)-(Var(Y)+(E(Y))2) =(9+22)-(4+02)=9

E(X+Y)*E(X-Y)=(2+0)*(2-0)=4

hence Cov((X+Y),(X-Y))=E((X+Y)*(X-Y))-E(X+Y)*E(X-Y)=9-4=5

therefore Corr((X+Y),(X-Y)) =Cov((X+Y),(X-Y))/sqrt(Var(X+Y)*Var(X-Y))

=5/sqrt(16*10)=5/sqrt(160)=0.3953